Recently, More Money 4U offered an annuity that pays 4.8% compounded monthly. If $1,453 is deposited into this annuity every month, how much is in the account after 7 years? How much of this is interest?

Type the amount in the account: $____

Type the amount of interest earned: $____

(Please square, label, or circle the answer)
Answer #1

Monthly rate of interest = 4.8%/12 = 0.4%

Number of months = 7*12 = 84 months

Future value of annuity = Periodic Amount*[{(1+r)^n - 1}/r]

Hence, amount after 7 years = 1453*[{(1+0.4%)^84 - 1}/0.004]

= $144,719.34

Interest amount = Total amount - Amount deposited

= 144,719.34 - 1453*84

= $22,667.34